OpenStudy (anonymous):

Which equation could have been used to create this function table? A. y = x + 9 B. y = x + 10 C. y = 9x D. y = 10x

OpenStudy (tiff9702):

what function table

OpenStudy (anonymous):

wheres the table

OpenStudy (anonymous):

Here is the table:

OpenStudy (tiff9702):

a

OpenStudy (anonymous):

Could you help me with a few more?

OpenStudy (tiff9702):

sure

OpenStudy (anonymous):

Which equation could have been used to create this function table? A. y = 3x B. y = x + 3 C. y = 4x D. y = x + 2

OpenStudy (tiff9702):

a is the answer again

OpenStudy (anonymous):

Victoria learns 2 songs each week in piano lessons. Which graph shows the number of songs Victoria will learn if she takes lessons for w weeks?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Sorry have to post them seperatly.

OpenStudy (anonymous):

why are there like 10 people here???

OpenStudy (anonymous):

I dunno. I have 1 more question left.

OpenStudy (tiff9702):

i am going with b

OpenStudy (anonymous):

Which graph shows the equation y = x + 2?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (tiff9702):

i think A but i am not sure

OpenStudy (anonymous):

Mary created a correct function table using the equation t = 8s Which table could be Mary's?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

B

OpenStudy (anonymous):

C

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Last one promise.

OpenStudy (tiff9702):

C

OpenStudy (anonymous):

Got it, will let you know if they were correct.

OpenStudy (tiff9702):

ok if you arer in K12's math yellow program i should be right (My brother is in it 2)

OpenStudy (anonymous):

First one is correct, Second one is too, Third one as well, the 4th and 5th were not. I got it 60% correct but i did not pass. :(

OpenStudy (tiff9702):

sorry

OpenStudy (anonymous):

Should i retry it?

OpenStudy (tiff9702):

yes (but also review)

OpenStudy (anonymous):

Ok, i will give answers just in case someone else needs some help/